OpenStudy (anonymous):

Solve the system of equations x-4y=5 x=3y-2 (_,_)

OpenStudy (anonymous):

@bibby help?

OpenStudy (bibby):

use substitution, solve for x using one of the equations and replace it in the other

OpenStudy (anonymous):

I'm confused.

OpenStudy (bibby):

never mind x=3y-2 x-4y=5 3y-2-4y=5

OpenStudy (anonymous):

Oh. But my answer is supposed to be an ordered pair, how do I know what it is?

OpenStudy (bibby):

solve fo for y...

OpenStudy (bibby):

then plug into one of the equations to find x

OpenStudy (anonymous):

Umm (-23,-7)?

OpenStudy (bibby):

show me how you're getting that \(3y-2-4y=5\) solve for y

OpenStudy (anonymous):

-y=-3?

OpenStudy (bibby):